Tony formed a business called Tone’s Big Screen TV Emporium Inc. during the month of September. Doors will open for business on October 1. The following occurred during September:

  1. Tony invested the following

    • Tony personally borrowed $10,000 from the bank. Tony then invested the funds in the business.

    • Used computer equip. $4,800, originally costing $6,000. Estimates 4 years of service remain.

    • Office equip. worth $2,400, originally costing $6,000. Estimates 5 years of service remain.

  2. Called Sony and ordered 10 big screen TVs at a cost of $500 each. Payment is due 30 days after receipt.

  3. Hired a salesperson who will be paid $500 per month plus commissions of 5% of what he sells. Salesperson

    will begin work on October 1.

  4. Signed a five-year lease agreement with rent due of $1,000 per month. Tone will move in the first of next

    month and pay first and last month’s rent.

  5. Tony received the 10 big screen TVs ordered from Sony.

  6. Called the advertising department of the local newspaper and arranged to run an opening day advertisement.

    Paid $1,000 for the ad which will run on October 1.

  7. Purchased new warehouse equipment for $8,000 in cash. The equipment will be useful for 5 years.

The following events occurred during the month of October:

  1. Borrowed $25,000 from the bank.

  2. Paid $2,000 on October 1st as required by the lease.

  3. Sold 4 TVs at $800 each for a $3,200 total. The cash was collected immediately.

  4. Paid Sony for the TVs ordered and received in September.

  5. Sold 4 TVs at $950 each - $3,800 total. The customer will pay for the TVs in November.

  6. Paid salesman his salary of $500 at the end of the month. Commissions will be paid next month.

  7. Customer signed sales contract for one TV at $850. The TV will be delivered next month. The customer

    made an advance payment of $400.

  8. Received the following bills for October services to be paid in November

    • $300 for October Utilities.

    • $50 for 2 months interest on Tone’s personal loan and $40 for one month on the company loan.

  9. Paid $500 in dividends to Tone in his capacity as a stockholder/owner.

(1) Record transactions (2) Prepare any necessary adjustments (3) Prepare financial statements.
Answer #1

Tonys Big Screen TV Emporium Inc. BALANCE SHEET FOR THE MONTH OF OCTOBER Current Assets Cash and Cash Equivalents Accounts R

Depreciation for the month of october
Computer Equipments 4800 4800/4=1200 1200/12=100
Office Equipments 2400 2400/5=480 480/12=40
Warehouse Equipments 8000 8000/5= 1600 1600/12=133.33

Tonys Big Screen TV Emporium Inc. Income Statement for the month october 7850 0 sales cost of goods sold Opening Inventories